OpenStudy (nuccioreggie):

The table below shows the number of marbles of different colors in a bag: Marble Experiment Color of Marbles Number of Marbles Red 1 White 5 Black 10 Ursula draws a marble from the bag randomly without looking. She then draws another marble from the bag without replacing the first one. Which expression shows the probability of drawing black marbles in both the trials? 10 over 16 added to 9 over 16 10 over 16 added to 9 over 15 10 over 16 multiplied by 9 over 16 10 over 16 multiplied by 9 over 15

OpenStudy (nuccioreggie):

@welshfella please help

OpenStudy (nuccioreggie):

thank you for coming really

OpenStudy (welshfella):

P(black first draw) = 10/16 p(black on 2nd draw) = 9/15 multiply these 2 probabolities

OpenStudy (nuccioreggie):

ok 1sec i am right now

OpenStudy (welshfella):

which one do you think it is?

OpenStudy (nuccioreggie):

3/6

OpenStudy (nuccioreggie):

oh oh oh sorry D

OpenStudy (welshfella):

no need to work it out yes thats right D

OpenStudy (nuccioreggie):

ok can you help me with a cople more please @welshfella

OpenStudy (welshfella):

gotta go right now sorry